Solve the following equation by using FOIL:
x^2-6x-17=0

Found 2 solutions by stanbon, rapaljer:
Answer by stanbon(75887)

Solve the following equation by using FOIL:
x^2-6x-17 = 0
-----
You cannot factor this by using inverse-Foil
---
FOIL itself is used to find products, not factors.
=================
Cheers,
Stan H.

Answer by rapaljer(4671)

This is called a Quadratic Equation, since it has an x^2 term. Since it's already equal to zero, in order to the next step is to FACTOR the trinomial. The trouble is, it doesn't factor. Are you sure you didn't copy the problem wrong?
x^2 - 6x - 17 = 0
(x- )(x+ ) = 0

My Email address is rapaljer@seminolestate.edu if you need to contact me!!

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus